Calculate the number of electrons constituting one coulomb of charge.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Here we will proceed by writing the formula for electrons constituting one coulomb of charge after that by using Milkman’s equation we will get our required answer. Complete Step-by-Step solution: Electron charge, (symbol e), fundamental physical constant expressing the naturally occurring unit of electron charge, equal to 1.602176634×10−19 coulomb. We need to calculate the number of electrons constituting one coulomb of charge. e=1.6×10−19C Total charge required 1 Coulomb. Therefore, q = 1C Hence using Milkman’s equation, we will get q=n×e Where, n= number of electrons constituting on charge n=qe -- (i) Hence, 1C charge is carried by 1C =n×1.6×10−19 C Substituting the value in equation (i) we get, ⇒n=11.6×10−19 =6.25×1018electrons. So, if 1A current flows through a conductor, then 6.25×1018 electrons pass per a second across the cross section of the conductor. Note: Whenever we come up with this type of question first we will write the formula for electrons constituting one coulomb of charge then with the help of formula we will derive the current flows through a conductor.